What is the minimum distance between a point on the curve y=e^x and a point on y=log x?

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Aditya,

its distance = sqrt ( (3-x)^2 + (2-y)^2 ) right?

So since y = x² + 2, you can substitute this into the above so that you have a function of only x, no y in it

d = sqrt ( (3-x)^2 + (2 - x² - 2)^2 )

Now, how do you minimize d? I assume you know how to do differential calculus because you should for this question

the first derivative is 1/2*(-6+2*x+4*x^3)/sqrt(9 - 6*x + x^2+ x^4) you set it to 0 and solve for x, You get 3 roots, one of which is real, the other are complex. Discard the complex roots, and take x = 1, which will satisfy the derivative = 0

Substitute x = 1 into d = sqrt ( (3-x)^2 + (2 - x² - 2)^2 ),

and you get d = sqrt(5) or 2.236067977

in fact x = 1, y = 3, is the point on the curve closest to the point (3, 2)

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert